# How General Are Measures of Choice Consistency? Evidence from Experimental and Scanner Data

### v1

@@ 5 Underlying Mechanisms
Our findings indicate that while consistency scores can be derived from choice data across various contexts, they do not always reflect a general ability to make decisions aligned with individual preferences. This resonates with existing literature highlighting the challenges in inferring true preferences from observable choices. For instance, Ariely et al. (2003) introduce the notion of coherent arbitrariness, suggesting that subjects are swayed by arbitrary anchors but respond consistently to significant changes in price, quantity, and quality. Similarly, Dean and Sautmann (2021) argue that intertemporal choices in experiments may not accurately represent true time preferences, as they can be affected by financial shocks and budget constraints. Furthermore, the frequently observed S-shaped probability-weighting function in prospect theory may arise from cognitive uncertainty, where individuals tend to default to cognitive shortcuts (Enke and Graeber, 2023), or from the complexities involved in evaluating lotteries (Oprea, 2024). Observable choices such as those in the Allais paradox and Ellsberg paradox may also reflect noise or errors (Nielsen and Rehbeck, 2022; McGranaghan et al., 2024) or cognitive challenges such as the failure of contingent thinking (Esponda and Vespa, 2024; Niederle and Vespa, 2023). Additionally, de Clippel et al. (2024) show that subjects, whose choices are consistent with utility maximization, often rely on heuristic decision rules that mimic the characteristics of utility functions rather than truly maximize utility. In a similar vein, consistency scores inferred from observable choices may be “as if” to some degree, because it is challenging to separate the influences of unobservable factors such as randomness, budget constraints, preferences, and heuristics which we elaborate below.
@@ 5.1 Randomness in Choice Behavior
We examine the randomness in the choice behaviors, which can be viewed as a measure of some unobservable factors. We apply a non-parametric one-sided test proposed by Cherchye et al. (2023) to examine whether a choice dataset is best explained by a random DM or an approximate utility maximizer. This test uses the permutation approach, which fixes the budget sets but permutes the relative consumption shares over different rounds for each DM, to operationalize the random behaviors. They show that, given a sufficient number of observed choices, the CCEI from an approximate utility maximizer has a probability close to 1 of exceeding CCEI generated by the permutation test. Accordingly, we compute the proportion of permuted CCEI scores, generated through 10,000 permutations, that surpass the observed CCEI. [^11] We follow the procedure in Cherchye et al. (2023) to abort the test when the proportion exceeds 0.2 for the first 1,000 permutations and report this value to speed up computation. If such proportion is smaller than the traditional significance level $\alpha=0.05$, we can regard the DM as an approximate utility maximizer.
@@
N $\alpha=0.01$ $\alpha=0.05$ $\alpha=0.1$
@@
We find that 77.06% participants in Experiment 1 can be classified as approximate utility maximizers, while this number is 22.65% when their scanner data are used. Furthermore, an individual classified as an approximate utility maximizer in the experiment does not have a significantly higher chance to be labeled as an approximate utility maximizer in the supermarket (22.5% vs. 23.1%, $p>0.1$, two-sample proportion test). We also examine the proportion in Experiment 2 and Scanner Dataset 2 and use alternative significance levels. Table 4 summarizes these proportions with the corresponding significance level, and shows that the results are robust. Our observations are comparable to those of Cherchye et al. (2023), where the proportions are 82% for social task in a two-goods setting in the experiment and 30% for real-life choices at the significance level of 0.05. Overall, this analysis shows that choices in scanner data are more random than those in the experiment, and the sources of randomness are distinct in these two settings.
@@ 5.2 Budget Constraints
One important difference across the two environments is the budget constraint. In the lab, participants are presented with well-defined budget lines using a simple interface. By contrast, the budget lines of consumers are not directly observable and are constructed by aggregating the choices in a given set of categories. In addition, consumption data often lack the power to detect GARP violations, because budget lines may not sufficiently cross (Blundell et al., 2003, 2008; Chambers and Echenique, 2016, Chapter 5). However, our scanner data do have the power to detect GARP violations. To take into account of the perceived budget constraints, we utilize two alternative approaches. The first approach is the Price Misperception Index (PMI) in de Clippel and Rozen (2023), which quantifies the minimal degree of price misperception that can rationalize choice data. The second approach is proposed by Deb et al. (2023) to examine the consistency with the generalized axiom of price preference (GAPP). This approach allows us to calculate deviations from consistent welfare predictions across different budgets, after accounting for one specific type of inattention to prices. [^12] Given $p^{1}=(1,2),x^{1}=(0,2);p^{2}=(2,1),x^{2}=(3,1)$, we can find the dataset satisfies GARP. However, it violates GAPP, because the participant can spend less to purchase bundle $x^{1}$ at period 2, thus being better off at price $p^{2}$ than at $p^{1}$. Similarly, she can also spend less to purchase bundle $x^{2}$ at period 1 and become better off at $p^{1}$ than at $p^{2}$. This leads to a contradiction in the preference over prices. GAPP is a necessary and sufficient condition for the dataset to be rationalized with $U(x,-px):\mathbb{R}_{+}^{K}\times\mathbb{R}_{-}\rightarrow\mathbb{R}$ (Theorem 1, Deb et al., 2023). According to Proposition 3 in Deb et al. (2023), since the expenditure in our experiment is fixed, their measure ($\vartheta^{*}$) is equivalent to CCEI. Panel A in Table 5 shows that the observed correlations are similar to those using CCEI. [^13] Additionally, we also conduct a few more robustness checks. First, we restrict our analyses to the most frequently purchased subcategory, i.e. pork and leaf vegetables, within each category and calculate the consistency scores accordingly. Second, we assume the price in each category is the same for all consumers in each month (Echenique et al., 2011; Dean and Martin, 2016). Specifically, we obtain the aggregate price index $P_{I}^{t}$ for category $I$ in month $t$: $P_{I}^{t}=\nicefrac{{\sum_{i\in I}e_{i}^{t}}}{{Q_{I}^{t}}}$, where $e_{i}^{t}$ represents the total expenditure of all consumers on item $i$ in category $I$ in month $t$, and $Q_{I}^{t}$ is the total quantity of items purchased in category $I$ in month $t$. In these analyses, all coefficients are close to our baseline findings.
Table 5: Correlations of Alternative Indices for Budgets Constraints and Heuristic Rules
Exp. 1 vs. Scanner Dataset 1 Exp. 2 Scanner Dataset 2
@@
  • Notes: / / /  represents that the magnitude of correlation coefficient is in [0,0.1)/[0.1,0.3)/[0.3,0.5)/[0.5,1]. Specific values are detailed in Table A8. -- denotes that the index is not applicable. Within Experiment 2, the three coefficients represent the correlations of consistency scores between risk and social tasks, risk and food tasks, and social and food tasks, respectively. In Scanner Dataset 2, the three coefficients across categories represent those between (Snacks,Meat) and (Veg.,Fruits), (Snacks,Veg.) and (Meat,Fruits), and (Snacks,Fruits) and (Meat,Veg.); those across time are correlations between (2018,2019) and (2020,2021), (2018,2020) and (2019,2021), and (2018,2021) and (2019,2020).
Constructing budget lines is inherently challenging due to various factors influencing both consumer behaviors and firm pricing strategies. On the demand side, consumers often respond to price changes in a coarse manner, leading to suboptimal purchasing decisions. For instance, they may exhibit left-digit bias, where they overemphasize the significance of changes in the leftmost digit of a price (List et al., 2023), and tend to evaluate price changes based on their memory of past prices, which can distort their perceptions of current value (Bordalo et al., 2020). On the supply side, firms frequently engage in strategic pricing and adjust their prices in ways that can manipulate consumer perceptions and purchasing behaviors. As a result, budget constraints are not exogenously given, but influenced by market dynamics. Addressing these complexities requires further research to accurately measure and understand these factors.
@@ 5.3 Formation of Preferences
Formation of preferences also differs across environments. Plott (1996) proposes the notion of preference discovery, whereby choice consistency with utility maximization can be understood as a discovery process in which subjects learn their own needs through a process of reflection and practice. This perspective highlights the dynamic nature of formation of preferences, where individuals refine their understanding of their needs and objectives over time. Complementing this view, Deming (2021) distinguishes between the decision-making ability of performing a routine task and a problem-solving task, and suggests that the formation of preferences may differ depending on the task characteristics. In this regard, DM as a consumer is more likely to do a routine task and may be subjective to some preference shocks related to various contextual factors. By contrast, when participants in the lab are presented with risky decisions and social decisions, they need to learn the abstract experimental rules, since these rules are not part of their daily experience. Therefore, choice consistency in the lab may be related to the ability to “discover” one’s preference in the new choice environment and “solve” the choice problem accordingly. Below we present some results testing the effect of some contextual factors.
@@
We examine whether consumers’ preferences on meat and vegetables change across time (years, seasons, days and hours) and other contexts (discounted vs. non-discounted purchases) in Scanner Dataset 1.
@@
Table 6 reports the results. By conducting pairwise comparisons among the four seasons, we observe that consumers’ food preferences in spring are notably different from other seasons, especially fall and winter. [^15] We also apply the method by Echenique et al. (2011). Specifically, we calculate the proportion of GARP violations across every two choices within specific pairwise combination, then take the average across consumers. We find the results are similar (Table A9 and A10). By estimating the CES functions across seasons, we further observe $\hat{\alpha}$ in spring is significantly lower than in other three seasons ($p<0.01$ for all comparisons, paired $t$-tests), indicating that consumers shift their preferences towards vegetables rather than meat in spring (Table A11). Two possible explanations exist. First, the majority of vegetables reach their optimal freshness in spring. Second, following the Chinese New Year—a significant temporal landmark marking the end of winter in China, individuals tend to set new year resolution, such as improving their diet or adopting healthier eating habits. Additionally, our analysis reveals distinct preferences across multiple dimensions: between meal and non-meal time, and between discounted and non-discounted purchases. Meanwhile, we cannot detect significant differences between the year 2020 and 2021 and between working and non-working days. Overall, these results suggest that there are contextual factors influencing choices in the supermarket setting.
@@
We analyze whether there is a learning effect in the experiment. We split 22 rounds of choices in Experiment 1 into the first half and the second half, and find that CCEI scores are lower for the first half of the risky decisions (0.967 vs. 0.981, $p<0.01$, paired $t$-test). Using the same method as mentioned above, the underlying preferences for the two halves are also different, contributing to the deviation from utility maximization. Moreover, given that we randomize the order of all three tasks in Experiment 2, we find that the Spearman’s rank correlation between the first and second tasks, and that between the first and third are both 0.284 ($p<0.01$), whereas that between the last two tasks is 0.350 ($p<0.01$). These findings suggest that the generalizability increases as participants gain more experience throughout the experiment.
@@
We further examine how individual characteristics affect consistency scores in these two settings. The independent variables include Raven’s IQ test score, shopping frequency, Big Five personality traits as well as demographic variables including gender, family income, education, age and family size. The dependent variables are CCEI scores (Column 1) and consistency with FOSD (Column 2) in Experiment 1, as well as CCEI scores in Scanner Dataset 1 (Column 3).
Table 7 presents OLS regression results. Our results show that the estimated coefficient of IQ is significantly positive for consistency with FOSD, which is similar to the observation from Cappelen et al. (2023). Moreover, consumers with more shopping experience have significantly higher scores of CCEI in the supermarket. This is in line with List (2003), showing that individuals with more trading experiences exhibit less endowment effect and behave closer to neoclassical predictions, and suggesting that consumers can learn to overcome the endowment effect. Given the maximum score of these measures is 1 and their distributions are skewed, we also apply a Tobit model and use relative rankings as dependent variables separately and show that the results remain robust (Table A12 and A13).
@@
Exp: CCEI Exp: FOSD Scanner: CCEI
(1) (2) (3)
IQ 0.003 0.013*** -0.000
(0.002) (0.004) (0.001)
Frequency 0.059 0.057 0.241***
(0.074) (0.155) (0.042)
Conscientiousness 0.002 0.003 -0.002
(0.002) (0.003) (0.001)
Extraversion $0$0.001 0.001
(0.002) (0.003) (0.001)
Agreeableness -0.003 -0.004 0.000
(0.002) (0.004) (0.001)
Openness -0.003* -0.005 0.002*
(0.002) (0.004) (0.001)
Neuroticism -0.004*** -0.006* -0.000
(0.002) (0.004) (0.001)
Female 0.000 0.027 -0.009*
(0.007) (0.017) (0.005)
Medium Income -0.004 0.013 -0.002
(0.008) (0.017) (0.005)
High income 0.012 0.043** -0.003
(0.009) (0.018) (0.005)
Medium education -0.004 -0.004 -0.003
(0.008) (0.016) (0.004)
High education 0.004 0.022 -0.005
(0.008) (0.017) (0.006)
Medium age 0.016 0.055* -0.004
(0.016) (0.033) (0.008)
Elder age 0.014 0.062* -0.005
(0.017) (0.035) (0.009)
Family size -0.002 -0.004 0.001
(0.003) (0.007) (0.002)
Simulated CCEI 0.406***
Constant 0.910*** 0.673*** 0.587***
(0.025) (0.054) (0.047)
N 1,055 1,055 1,055
R2 0.023 0.034 0.084

### v2

@@ 5 Underlying Mechanisms
Our findings indicate that, while consistency scores can be derived from choice data across various contexts, they do not always reflect a general ability to make decisions aligned with individual preferences. This resonates with the existing literature that highlights the challenges in inferring true preferences from observable choices. For example, Ariely et al. (2003) introduce the notion of coherent arbitrariness, suggesting that subjects are swayed by arbitrary anchors but respond consistently to significant changes in price, quantity, and quality. Similarly, Dean and Sautmann (2021) argue that intertemporal choices in experiments may not accurately represent true time preferences, as they can be affected by financial shocks and budget constraints. Furthermore, the frequently observed S-shaped probability weighting function in prospect theory may arise from cognitive uncertainty, where individuals tend to default to cognitive shortcuts (Enke and Graeber, 2023), or from the complexities involved in evaluating lotteries (Oprea, 2024). Observable choices such as those in the Allais paradox and the Ellsberg paradox may also reflect noise or errors (Nielsen and Rehbeck, 2022; McGranaghan et al., 2024) or cognitive challenges such as the failure of contingent thinking (Esponda and Vespa, 2024; Niederle and Vespa, 2023). Moreover, de Clippel et al. (2024) show that subjects whose choices are consistent with utility maximization often rely on heuristic decision rules that mimic the characteristics of utility functions rather than truly maximize utility. Similarly, the consistency scores derived from observable choices might appear “as if” they are influenced by unobservable factors such as randomness, budget constraints, preferences, and heuristics, making it challenging to isolate these effects. We will delve into these factors to assess their role in the observed lack of correlation between choice consistency in supermarket and experimental data. This analysis will help shed light on whether the lack of correlation is due to measurement errors or the multidimensional nature of choice consistency across different contexts.
@@ 5.1 Randomness in Choice Behavior
We examine the randomness of choice behaviors, which can be viewed as a measure of some unobservable factors. We apply a non-parametric one-sided test proposed by Cherchye et al. (2025) to examine whether a choice dataset is best explained by a random DM or an approximate utility maximizer. This test uses the permutation approach, which fixes the budget sets but permutes the relative consumption shares over different rounds for each DM, to operationalize the random behaviors. They show that, given a sufficient number of observed choices, the CCEI of an approximate utility maximizer has a probability close to 1 of exceeding the CCEI generated by the permutation test. Consequently, we calculate the proportion of permuted CCEI scores, generated through 10,000 permutations, that surpass the observed CCEI. [^7] We follow the procedure in Cherchye et al. (2025) to abort the test when the proportion exceeds 0.2 for the first 1,000 permutations and report this value to speed up the computation. If such a proportion is smaller than the traditional significance level $\alpha_{\text{sig}}=0.05$, we can regard the DM as an approximate utility maximizer.
@@
N $\alpha_{\text{sig}}=0.01$ $\alpha_{\text{sig}}=0.05$ $\alpha_{\text{sig}}=0.1$
@@
We find that 77.06% participants in Experiment 1 can be classified as approximate utility maximizers, while this number is 22.65% when their scanner data are used. Furthermore, an individual classified as an approximate utility maximizer in the experiment does not have a significantly higher chance to be labeled as an approximate utility maximizer in the supermarket (22.5% vs. 23.1%, $p>0.1$, two-sample proportion test). We also examine the proportion in Experiment 2 and Scanner Dataset 2 and use alternative significance levels. Table 4 summarizes these proportions with the corresponding significance level and shows that the results are robust. Our observations are comparable to those of Cherchye et al. (2025), where the proportions are 82% for the social task in a two-goods setting in the experiment and 30% for real-life choices at the significance level of 0.05. Overall, this analysis shows that the choices in the scanner data are more likely from random DMs than those in the experiment, and the sources of randomness are distinct in these two settings.
@@ 5.2 Budget Constraints
An important difference between the two environments is the budget constraint. In the experiment, participants are presented with well-defined budget lines using a simple interface. In contrast, the budget lines of consumers are not directly observable and are constructed by aggregating the choices in a given set of categories. In addition, consumption data often lack the power to detect GARP violations, because budget lines may not cross sufficiently (Blundell et al., 2003, 2008; Chambers and Echenique, 2016, Chapter 5). However, our scanner data do have the power to detect GARP violations. To further take into account perceived budget constraints, we utilize two alternative approaches. The first approach is the Price Misperception Index (PMI) in de Clippel and Rozen (2023), which quantifies the minimal degree of price misperception that can rationalize choice data. The second approach is proposed by Deb et al. (2023), which examines the consistency with the Generalized Axiom of Price Preference (GAPP). This approach allows us to calculate deviations from consistent welfare predictions across different budgets, after accounting for one specific type of inattention to prices. [^8] Given $p^{1}=(1,2),x^{1}=(0,2);p^{2}=(2,1),x^{2}=(3,1)$, we can find the dataset satisfies GARP. However, it violates GAPP, because the participant can spend less to purchase bundle $x^{1}$ in period 2, thus being better off at price $p^{2}$ than at $p^{1}$. Similarly, she can also spend less to purchase the bundle $x^{2}$ in period 1 and become better off at $p^{1}$ than at $p^{2}$. This leads to a contradiction in the preference for prices. GAPP is a necessary and sufficient condition for the dataset to be rationalized with $U(x,-px):\mathbb{R}_{+}^{K}\times\mathbb{R}_{-}\rightarrow\mathbb{R}$ (Theorem 1, Deb et al., 2023). According to Proposition 3 in Deb et al. (2023), since the expenditure in our experiment is fixed, their measure ($\vartheta^{*}$) is equivalent to CCEI, but this equivalence does not hold in the supermarket setting. Panel A in Table 5 shows that the observed correlations are similar to those using CCEI.
Table 5: Correlations of Alternative Indices for Budget Constraints and Heuristic Rules
Exp. 1 vs. Scanner
Dataset 1
Exp. 2 Scanner Dataset 2
@@
  • Notes: / / / represents that the magnitude of correlation coefficient is in [0,0.1)/[0.1,0.3)/[0.3,0.5)/[0.5,1]. Specific values are detailed in Table A8. - denotes that the index is not applicable. Within Experiment 2, the three coefficients represent the correlations of consistency scores between risk and social tasks, risk and food tasks, and social and food tasks, respectively. In Scanner Dataset 2, the three coefficients across categories represent those between (Snacks,Meat) and (Veg.,Fruits), (Snacks,Veg.) and (Meat,Fruits), and (Snacks,Fruits) and (Meat,Veg.); those across time are correlations between (2018,2019) and (2020,2021), (2018,2020) and (2019,2021), and (2018,2021) and (2019,2020).
The construction of budget lines is inherently challenging due to various factors that influence both consumer behavior and firm pricing strategies. On the demand side, consumers often respond to price changes in a coarse manner, leading to suboptimal purchasing decisions. For example, they may exhibit left-digit bias, where they overemphasize the significance of changes in the leftmost digit of a price (List et al., 2023), and tend to evaluate price changes based on their memory of past prices, which can distort their perceptions of current value (Bordalo et al., 2020). On the supply side, firms frequently engage in strategic pricing and adjust their prices in ways that can manipulate consumer perceptions and purchasing behaviors. As a result, budget constraints are not exogenously given, but influenced by market dynamics. Addressing these complexities requires further research to accurately measure and understand these factors.
@@ 5.3 Formation of Preferences
The formation of preferences also differs between environments. Plott (1996) proposes the notion of preference discovery, in which the choice consistency with utility maximization can be understood as a discovery process in which subjects learn their own needs through a reflection and practice process. This perspective highlights the dynamic nature of the formation of preferences, where individuals refine their understanding of their needs and objectives over time. Complementing this view, Cerigioni (2021), Deming (2021) and Ilut and Valchev (2023) distinguish between the decision-making ability to perform a routine task and a problem-solving task, and suggests that the formation of preferences may differ depending on the characteristics of the task. In this regard, DM as a consumer is more likely to do a routine task and may be subjective to some preference shocks related to various contextual factors. In contrast, when participants in the experiment face risky decisions and social decisions, they need to learn the abstract experimental rules, as these rules are not part of their daily experience. Therefore, the choice consistency in the experiment may be related to the ability to “discover” one’s preference in the new choice environment and “solve” the choice problem accordingly. In the following, we present the results that test the effect of some contextual factors.
@@
We examine whether consumers’ preferences on meat and vegetables change over time (years, seasons, days, and hours) and other contexts (discounted vs. non-discounted purchases) in Scanner Dataset 1.
@@
Table 6 reports the results. By conducting pairwise comparisons among the four seasons, we observe that consumers’ food preferences in spring are notably different from other seasons, especially fall and winter. [^10] We also apply the method by Echenique et al. (2011). Specifically, we calculate the proportion of GARP violations across every two choices within a specific pairwise combination, then take the average across consumers. We find that the results are similar (Tables A9 and A10). By estimating the CES functions between seasons, we further observe that $\hat{\alpha}$ in spring is significantly lower than in other three seasons ($p<0.01$ for all comparisons, paired $t$-tests), indicating that consumers change their preferences towards vegetables rather than meat in spring (Table A11). There are two possible explanations. First, most vegetables reach their optimal freshness in spring. Second, following Chinese New Year—a significant temporal landmark that marks the end of winter in China, people tend to set a resolution for the new year, such as improving their diet or adopting healthier eating habits. Additionally, our analysis reveals distinct preferences across multiple dimensions: between meal and non-meal time, and between discounted and non-discounted purchases. Meanwhile, we cannot detect significant differences between the year 2019 and 2020 and between working and non-working days. In general, these results suggest that there are contextual factors that change preferences in the supermarket setting.
@@
[…]
@@
We further examine how individual characteristics affect consistency scores in these two settings in Table 7. Independent variables include Raven’s IQ test score, shopping frequency, Big Five personality traits, as well as demographic variables that include gender, family income, education, age, and family size. The dependent variables are CCEI scores (Column 1) and consistency with FOSD (Column 2) in Experiment 1, as well as the CCEI scores in Scanner Dataset 1 (Column 3).
Our results show that the estimated coefficient of IQ is significantly positive for the consistency with FOSD, which is similar to the observation from Cappelen et al. (2023). In addition, consumers with more shopping experience have significantly higher CCEI scores in the supermarket. This is in line with List (2003), showing that individuals with more trading experiences exhibit less endowment effect and behave more closely to neoclassical predictions, and suggesting that consumers can learn to overcome the endowment effect. Given that the maximum score for these measures is 1 and their distributions are skewed, we also apply a Tobit model and use relative rankings as dependent variables separately and show that the results remain robust (Tables A12 and A13).
Additionally, Column 4 builds on Column 3 by adding the CCEI score in Experiment 1 as an independent variable. The coefficient on this variable is statistically insignificant. According to the Frisch-Waugh-Lovell theorem (Frisch and Waugh, 1933; Lovell, 1963), this regression result is equivalent to showing that the residuals from regressing CCEI in Experiment 1 on individual characteristics are not significantly correlated with the residuals from regressing CCEI in Scanner Dataset 1 on the same set of characteristics. That is, Result 1 continues to hold after accounting for individual characteristics in each setting.
@@
Exp: CCEI Exp: FOSD Scanner: CCEI Scanner: CCEI
(1) (2) (3) (4)
IQ 0.003 0.013*** -0.000 -0.000
(0.002) (0.004) (0.001) (0.001)
Frequency 0.059 0.057 0.241*** 0.241***
(0.074) (0.155) (0.042) (0.042)
Conscientiousness 0.002 0.003 -0.002 -0.002
(0.002) (0.003) (0.001) (0.001)
Extraversion 0.000 -0.001 0.001 0.001
(0.002) (0.003) (0.001) (0.001)
Agreeableness -0.003 -0.004 0.000 0.000
(0.002) (0.004) (0.001) (0.001)
Openness -0.003* -0.005 0.002* 0.002*
(0.002) (0.004) (0.001) (0.001)
Neuroticism -0.004*** -0.006* -0.000 -0.000
(0.002) (0.004) (0.001) (0.001)
Female 0.000 0.027 -0.009* -0.009*
(0.007) (0.017) (0.005) (0.005)
Medium Income -0.004 0.013 -0.002 -0.002
(0.008) (0.017) (0.005) (0.005)
High income 0.012 0.043** -0.003 -0.003
(0.009) (0.018) (0.005) (0.005)
Medium education -0.004 -0.004 -0.003 -0.003
(0.008) (0.016) (0.004) (0.004)
High education 0.004 0.022 -0.005 -0.005
(0.008) (0.017) (0.006) (0.006)
Medium age 0.016 0.055* -0.004 -0.004
(0.016) (0.033) (0.008) (0.008)
Elder age 0.014 0.062* -0.005 -0.005
(0.017) (0.035) (0.009) (0.009)
Family size -0.002 -0.004 0.001 0.001
(0.003) (0.007) (0.002) (0.002)
Simulated CCEI 0.406*** 0.406***
(0.052) (0.052)
Exp: CCEI -0.005
Constant 0.910*** 0.673*** 0.587*** 0.592***
(0.025) (0.054) (0.047) (0.051)
N 1,055 1,055 1,055 1,055
R2 0.023 0.034 0.084 0.084